# StyleStudio: Text-Driven Style Transfer with Selective Control of Style Elements

### v1

@@ 2 Related Work
Style Transfer. Style transfer refers to the process of applying the style of a specific reference image to a target content image. Depending on the source of the content image, this task can be broadly categorized into two main approaches: image-driven style transfer and text-driven style transfer. Both approaches share the common goal of extracting and applying more refined and precise styles, which is an ongoing pursuit in the field. Image-driven style transfer faces the challenge of retaining the essential features of the content image in the generated result. For instance, InST [39], a textual inversion-based method, employs stochastic inversion to preserve critical details in the content image. StyleID [2] propose query preservation to maintain original content. Similarly, InjectFusion [13] enhances style transfer outcomes by effectively blending features from the h-space [16]. Recently, text-driven style transfer has focused on addressing content leakage, where the application of style could disrupt or obscure important content features. B-LoRA [7] finds that jointly learning the LoRA weights of two specific blocks could achieves style-content separation. DEADiff [23] improves model performance by training joint image-text cross-attention mechanisms, leading to enhanced generation quality. StyleAlign [35] employs a strategy of swapping the query and key components within self-attention mechanisms to achieve style-aligned image generations, effectively aligning the stylistic attributes of the output with the desired style input.
Adapter-Based Stylization Methods. IP-Adapter [37] introduces a Decoupled Cross Attention mechanism to inject image information into the final generated output. Benefiting from training on large-scale datasets, it is capable of performing zero-shot style transfer tasks. Additionally, by employing weight tuning, IP-Adapter [37] can effectively alleviate content leakage issues. Techniques such as InstantStyle [33] mitigate content leakage by selecting appropriate layers in Stable Diffusion’s UNet [26, 22, 27] for injecting style image features, which avoiding the need for complex manual parameter tuning. StyleShot [9] extracts style information at multiple levels of the image encoder to enable fine-grained control over the style application, allowing for more precise adjustments to the generated image’s stylistic characteristics. CSGO [36] advances the field by utilizing a carefully prepared style dataset to train an adapter-based model that seamlessly decouples and injects content and style features.
@@ 3.1 Preliminaries
Style Transfer with an Adapter. Recent advancements in style transfer have explored the use of adapter-based methods to inject style-specific features into text-to-image models. These adapter networks are designed to modify pre-trained models to accommodate new styles inputs. One notable approach is the IP-Adapter [37], which leverages a dual cross-attention mechanism to integrate both text and image conditions. The CSGO [36] shares the same architecture for text-driven style transfer, with the key distinction being that it is specifically trained on a style transfer dataset, enabling it to better capture and transfer stylistic features. In this design, one cross-attention layer extracts features from the text prompt, while the other extracts features from the image. The two sets of features are then combined through a weighted sum, allowing the model to generate images that adhere to both the stylistic attributes specified by the text and the visual characteristics of the image. This technique provides a simple yet effective method for adapting existing models to perform style transfer without the need for extensive retraining or additional model parameters. This combination process can be formalized as:
@@ 3.2 Cross-Modal AdaIN of Text and Style Conditioning
In typical text-driven style transfer tasks, text conditioning primarily serves to define the content, dictating the semantic structure of the output, while image conditioning predominantly encodes the stylistic features, such as texture and color palette. However, directly combining these two conditioning inputs through weighted summation, as in existing methods [37, 33], forces them to assume similar roles in the fusion process. This can create conflicts, particularly when the text and image provide divergent information about the content and style. For example, a text prompt may describe a scene in one way, while the image may impose a stylistic choice that conflicts with this description, leading to suboptimal results. To address this challenge, we revisit Adaptive Instance Normalization (AdaIN) [12], a widely recognized technique in style transfer. AdaIN operates by normalizing the content input $x$ based on the statistical properties (mean and standard deviation) of the style input $y$, integrating style characteristics while preserving the essential structure of the content. The AdaIN process is defined as follows:
@@
As shown in  Fig. 5, building on AdaIN, we develop a cross-modal AdaIN mechanism that integrates text and style conditioning in a way that respects their distinct roles. The Adapter-Based methods [33, 36] employed a weighted sum approach for feature fusion, ensuring that feature maps operated within the same embedding space. We discovered that AdaIN could directly replace the weighted sum strategy. This substitution enables effective feature fusion without the need for additional training, making it particularly advantageous when integrated with Adapter-Based methods [33, 36]. Specifically, we first leverage the cross-attention layers within the U-Net architecture to query both the style and text features based on the U-Net feature map. This step results in two separate grid feature maps—one for the style and one for the text—that share the same spatial resolution. Next, we apply AdaIN between two feature maps, where the feature maps queried by text conditions are normalized by style feature maps. Finally, the normalized text feature maps are fused into the raw U-Net features by simple addition. This fusion is constructed in a residual design, similar to the approach used in the raw cross-attention layers. Mathematically, this adaptive fusion process can be expressed as follows:
$f_{\text{style}}$ $=A(Q,K_{\text{style}},V_{\text{style}}),$ (5)
$f_{\text{text}}$ $=A(Q,K_{\text{text}},V_{\text{text}}),$
@@
Figure 5: The illustration of our proposed cross-modal AdaIN.
@@ 3.3 Layout Stablization with Teacher Model
Figure 6: Pipeline for Layout Stabilization in Text-Driven Style Transfer Using a Teacher Model. During the denoising generation process, the Teacher Model (e.g., SDXL [22]) ensures layout stability, guiding the style transfer model to preserve the structure while applying the desired stylistic transformation.
@@
In the context of Text-Driven Style Transfer, maintaining a stable layout is crucial for ensuring that the generated image accurately reflects the structure and composition described by the textual prompt. Specifically, the layout refers to the spatial arrangement of objects, elements, and scene composition in the image. Without stable layout alignment, these elements may become misaligned, resulting in images where the content is distorted, or the intended focus, perspective, and balance outlined in the prompt are lost. To address this, we propose a method that stabilizes the layout by selectively replacing certain Self-Attention AttnMaps [32, 26, 22] in the stylized image with those from the original diffusion model. This selective replacement helps preserve the spatial relationships and arrangement of key features in the image, ensuring that the core layout remains consistent throughout the denoising process. By doing so, we retain the structural coherence of the original image while still applying the desired stylistic transformation, leading to a more coherent and faithful alignment with the textual prompt. The following figure,  Fig. 6, provides a visual overview of this process, showcasing the integration of layout stabilization within the generative framework.
@@
Unlike conventional image editing approaches that replace Self-Attention AttnMaps across all timesteps, our method selectively replaces AttnMaps only during the initial denoising stages. This choice is based on the observation that core layout features stabilize in the early denoising steps, while later stages primarily refine stylistic details. Applying full-timestep replacement would risk excessive loss of stylistic detail, as structural information would override key stylistic attributes, limiting the effectiveness of style transfer.
@@
3.4 Style-Based CFG
Classifier-Free Guidance Mechanism. In standard CFG, the model generates outputs conditioned on a given text prompt, as well as outputs generated without any conditioning (i.e., the unconditional model output). The final output is a weighted combination of these two, where the conditional output steers the generation process in alignment with the prompt, while the unconditional output acts as a form of negative conditioning, helping to prevent the model from producing undesirable features. The unconditional output in the Style-Based CFG can be replaced with conditioning on a negative prompt, such as ”blur” or ”artifact,” to actively discourage the generation of undesirable features. Therefore, the CFG [11] mechanism can be formalized as:
Style-Based Classifier-Free Guidance. In the context of style transfer, we extend this CFG mechanism to address the challenge of style ambiguity in images. Specifically, we introduce the concept of a negative style image that retains the overall content of the reference image but excludes the target style element. This negative image serves as a counterpart to the original style image and helps the model focus on transferring only the desired style component. To generate the negative style image, we utilize a layout-controlled generation model, such as ControlNet, which allows us to create an image $z_{t}^{\text{neg}}$ that preserves the structural features of the reference image but omits the target style. This negative image acts similarly to a negative prompt in text-guided CFG, effectively guiding the model towards emphasizing the desired style while avoiding the generation of undesired style elements. The SCFG mechanism operates as follows: 1) Generate a Negative Image: Using a layout-controlled model, create a negative sample image $z_{t}^{\text{neg}}$ that preserves the image’s structural elements but omits the target style. 2) Apply SCFG to Guide Generation: Formulate the Style-Based CFG by adjusting the balance between the target style image $z_{t}^{\text{style}}$ and the negative style image $z_{t}^{\text{neg}}$. This balance is controlled by a weight factor $w$, which determines the influence of each image in the generation process. The noise prediction is modified as follows:
$\begin{split}\hat{\epsilon}_{\theta}(z_{t},t,y)&=(1+w)\cdot\epsilon_{\theta}(z_{t},y_{\text{cond}},y^{s}_{\text{cond}})\\ &-w\cdot\epsilon_{\theta}(z_{t},y_{neg},y_{neg}^{s}),\end{split}$ (7)
By integrating SCFG, our method enhances the generation process by isolating specific style components, filtering out extraneous, and thus focusing style transfer precisely on the desired feature. This approach mitigates the risk of overfitting to irrelevant style components, allowing the model to perform effective style transfer in complex scenarios involving multiple stylistic elements.
4 Evaluation and Experiments
Figure 8: Impact of Style-Based CFG. From left to right, the prompts are: “A car glides along a street lined with trees shedding their colorful, orange and red leaves”, “A car drives down a sunny street”, “A cat”, and “A blue car”. The proposed Style-Based CFG successfully eliminates unintended style elements such as “snow” and “golden leaves”, while text-based CFG fails to address these spurious attributes.
Implementation details. We have implemented our method on top of the latest Adapter-Based Style Transfer approach, named CSGO [36]. To ensure fairness in comparison and mitigate the influence of random initialization, we fixed the initial noise for all methods, as the initial noise can significantly impact the final results [10]. Additionally, we set the random seed to 42 for reproducibility, used 50 inference steps, and applied a uniform guidance scale of 5 across all methods. In the qualitative and quantitative comparison experiments, for the implementation involving the Teacher Model, its participation was specifically limited to the first 20 steps. All experiments were conducted on a single NVIDIA GTX-4090 GPU.
Metric SDXL-based Methods SD15-based Methods Ours
infer Time $(s)$ 6 6 9 48 4 3 2 17
[…]
Cross-Modal AdaIN Teacher Model Text Alignment $\uparrow$
0.216
0.223 (+3.2%)
0.228 (+5.5%)
0.235 (+8.7%)
4.1 Comparison with State-of-the-Arts
Qualitative Comparisons. Fig. 7 illustrates the comparison results with the state-of-the-art methods. Our method effectively addresses the issue of style overfitting, which is a greater risk for methods like CSGO [36], styleshot [9] that are specifically trained on style data. While methods such as CSGO [36], InstantStyle [33], and DEADiff [23] excel in mitigating content leakage, they still struggle with prompt alignment, often failing to fully capture the intended subject matter or details specified in the prompts. On the other hand, StyleShot [9] and StyleCrafter [18] can suffer from content leakage, leading to generated images that do not accurately reflect the input prompts. Our method prioritizes text alignment by accurately capturing and reflecting the key stylistic attributes specified in the prompt, such as color. Additionally, it ensures a stable layout in the generated images, maintaining structural integrity while achieving high similarity to the target styles. More detailed results and explanations could be found in Appendix D.
4.2 Style-Based CFG
4.3 Ablation Study
Teacher Model. In Fig. 9, shows a visualization of cross-attention maps for the word “apple” in the prompt “A red apple” across different model configurations. The first row represents the baseline model, where the attention map fails to effectively focus on the core concept of “apple” from the text prompt. In the second row, the addition of cross-modal AdaIN provides some improvement, but the attention is still diffuse and lacks precise focus. In the third row, with the Teacher Model added, the attention map becomes more concentrated on the target concept, demonstrating a clear improvement in alignment with the prompt. This comparison indicates that it is the Teacher Model that enables the model to better capture and emphasize key elements from the text prompt, resulting in improved attention quality and image generation accuracy. In Fig. 10, this experiment demonstrates the importance of selecting an appropriate denoising timestep to stop the involvement of the Teacher Model. If the Teacher Model’s influence is removed too early, as seen at lower timesteps, issues with layout stability persist, resulting in compositions that lack coherence and include multiple instances of the target object (“apple”). However, if the Teacher Model is involved throughout the full denoising process (e.g., at timestep 50), there is a noticeable loss of style in the generated image. This suggests that while the Teacher Model is crucial for achieving layout stability, an extended involvement can dilute the style elements, making it essential to find a balanced point to discontinue its influence for optimal results. More detailed results and explanations could be found in Appendix C.
5 Conclusion
In conclusion, existing text-driven style transfer faces key issues such as style overfitting and layout instability, which limit the adaptability and coherence of generated images. To address these challenges, we proposed three methods: cross-modal AdaIN for harmonizing style and text features, style-based classifier-free guidance (SCFG) for selective control of stylistic elements, and a Teacher Model to enhance layout stability. Experimental results confirm that our approach effectively mitigates these issues, leading to improved alignment, stability, and control in style transfer, making it a versatile and robust solution for text-to-image synthesis tasks.
Appendix A Implementation Details

### v2

@@ 2 Related Work
Stylized Image Generation. Style transfer applies the style of a reference image to a target content image and can be broadly categorized into image-driven and text-driven approaches. Image-driven methods, such as InST [40], StyleID [2], and InjectFusion [14], focus on preserving content while injecting style, employing techniques like stochastic inversion, query preservation, and feature blending in the h-space [17]. Text-driven methods aim to mitigate content leakage, where excessive style application distorts content features. B-LoRA [7] achieves style-content separation via LoRA weight optimization, while DEADiff [24] enhances text-image consistency through joint cross-attention learning. StyleAlign [36] and Visual Style Prompt [13] modify self-attention mechanisms by swapping either query-key (Q-K) or key-value (K-V) pairs, ensuring style alignment. Adapter-based methods provide an efficient alternative by injecting style features into pre-trained diffusion models. IP-Adapter [38] leverages a decoupled cross-attention mechanism for zero-shot style transfer and mitigates content leakage via weight tuning. InstantStyle [34] improves stylization by selectively injecting style features into Stable Diffusion’s UNet [27, 23, 28]. StyleShot [9] extracts multi-level style features for fine-grained control, while CSGO [37] enhances adapter-based stylization by training on a curated style dataset for effective style-content decoupling.
@@ 3 Method
Figure 4: The illustration of our proposed Cross-Modal AdaIN, Teacher Model, Style-Based CFG.
@@ 3.1 Preliminaries
Style Transfer with an Adapter. Recent advancements in style transfer have explored the use of adapter-based methods to inject style-specific features into text-to-image models. These adapter networks are designed to modify pre-trained models to accommodate new styles inputs. One notable approach is the IP-Adapter [38], which leverages a dual cross-attention mechanism to integrate both text and image conditions. The CSGO [37] shares the same architecture for text-driven style transfer, with the key distinction being that it is specifically trained on a style transfer dataset, enabling it to better capture and transfer stylistic features. In this design, one cross-attention layer extracts features from the text prompt, while the other extracts features from the image. The two sets of features are then combined through a weighted sum, allowing the model to generate images that adhere to both the stylistic attributes specified by the text and the visual characteristics of the image. This combination process can be formalized as:
@@ 3.2 Cross-Modal AdaIN of Text and Style Conditioning
In typical text-driven style transfer tasks, text conditioning primarily serves to define the content, dictating the semantic structure of the output, while image conditioning predominantly encodes the stylistic features, such as texture and color palette. However, directly combining these two conditioning inputs through weighted summation, as in existing methods [38, 34, 37, 9], forces them to assume similar roles in the fusion process. This can create conflicts, particularly when the text and image provide divergent information about the content and style. For example, a text prompt may describe a scene in one way, while the image may impose a stylistic choice that conflicts with this description, leading to suboptimal results. To address this challenge, we revisit Adaptive Instance Normalization (AdaIN) [12], a widely recognized technique in style transfer. AdaIN operates by normalizing the content input $x$ based on the statistical properties (mean and standard deviation) of the style input $y$, integrating style characteristics while preserving the essential structure of the content. The AdaIN process is defined as follows:
@@
As shown in  Fig. 4(a), building on AdaIN, we develop a Cross-Modal AdaIN mechanism that integrates text and style conditioning in a way that respects their distinct roles. The Adapter-Based methods [38, 37] employed a weighted sum approach for feature fusion, ensuring that feature maps operated within the same embedding space. We discovered that AdaIN could directly replace the weighted sum strategy. This substitution enables effective feature fusion without the need for additional training, making it particularly advantageous when integrated with Adapter-Based methods [34, 37]. Specifically, we first leverage the cross-attention layers within the U-Net architecture to query both the style and text features based on the U-Net feature map. This step results in two separate grid feature maps—one for the style and one for the text—that share the same spatial resolution. Next, we apply AdaIN between two feature maps, where the feature maps queried by text conditions are normalized by style feature maps. Finally, the normalized text feature maps are fused into the raw U-Net features by simple addition. This fusion is constructed in a residual design, similar to the approach used in the raw cross-attention layers. Mathematically, this adaptive fusion process can be expressed as follows:
$f_{\text{style}}=A(Q,$ $K_{\text{style}},V_{\text{style}}),f_{\text{text}}=A(Q,K_{\text{text}},V_{\text{text}}),$ (5)
@@ 3.3 Layout Stablization with Teacher Model
In the context of Text-Driven Style Transfer, maintaining a stable layout is crucial for ensuring that the generated image accurately reflects the structure and composition described by the textual prompt. Specifically, layout refers to the spatial arrangement of objects, elements, and the overall scene composition in the image. Without stable layout alignment, these elements may become misaligned, resulting in images where the content is distorted, or the intended focus, perspective, and balance outlined in the prompt are lost. To address this, we propose a method that stabilizes the layout by selectively replacing certain Self-Attention AttnMaps [33, 27, 23] in the stylized image with those from the original diffusion model. This selective replacement helps preserve the spatial relationships and arrangement of key features in the image, ensuring that the core layout remains consistent throughout the denoising process. By doing so, we retain the structural coherence of the original image while still applying the desired stylistic transformation, leading to a more coherent and faithful alignment with the textual prompt. In Fig. 4(b), provides a visual overview of this process, showcasing the integration of layout stabilization within the generative framework.
@@
Unlike conventional image editing approaches [32] that replace Self-Attention maps (AttnMaps) across all timesteps and decoder layers, our method selectively replaces AttnMaps only during the initial denoising timesteps but across all layers of the UNet. Applying full-timestep replacement, as conventional methods do, risks excessive loss of stylistic details, limiting the effectiveness of style transfer.
3.4 Style-Based CFG
Classifier-Free Guidance Mechanism. In standard CFG, the model generates outputs conditioned on a given text prompt, as well as outputs generated without any conditioning (i.e., the unconditional model output). The final output is a weighted combination of these two, where the conditional output steers the generation process in alignment with the prompt, while the unconditional output acts as a form of negative conditioning, helping to prevent the model from producing undesirable features. The unconditional output in the Style-Based CFG can be replaced with conditioning on a negative prompt, such as ”blur” or ”artifact” to actively discourage the generation of undesirable features. Therefore, the CFG [11] mechanism can be formalized as:
Style-Based Classifier-Free Guidance. In the context of style transfer, we extend this CFG mechanism to address the challenge of style ambiguity in images. Specifically, we introduce the concept of a negative style image that retains the overall content of the reference image but excludes the target style element. This negative image serves as a counterpart to the original style image and helps the model focus on transferring only the desired style component. To generate the negative style image, we use a layout-controlled generation model, such as ControlNet, which allows us to create an image $z_{t}^{\text{neg}}$ that preserves the structural features of the reference image but omits the target style. This negative image functions similarly to a negative prompt in text-guided CFG, effectively guiding the model to emphasize the desired style while suppressing undesired style elements. As show in Fig. 4(c), the SCFG mechanism operates as follows: 1) Generate a Negative Image: Using ControlNet, generate a negative sample image $z_{t}^{\text{neg}}$ that retains the structural elements of the image while omitting the target style. 2) Apply SCFG to Guide Generation: Formulate the Style-Based CFG by modulating the balance between the target style image $z_{t}^{\text{style}}$ and the negative style image $z_{t}^{\text{neg}}$. This balance is governed by a weight factor $w$, which determines the contribution of each image during generation. The noise prediction is modified as follows:
$\begin{split}\hat{\epsilon}_{\theta}(z_{t},t,y)&=(1+w)\cdot\epsilon_{\theta}(z_{t},y^{text}_{\text{cond}},y^{style}_{\text{cond}})\\ &-w\cdot\epsilon_{\theta}(z_{t},y^{text}_{neg},y_{neg}^{style}),\end{split}$ (7)
By integrating SCFG, our method refines the generation process by isolating specific style components, filtering out extraneous ones, and thereby focusing style transfer on the desired features. This approach reduces the risk of overfitting to irrelevant style components, allowing the model to perform effective style transfer in complex scenarios with multiple stylistic elements.
4 Evaluation and Experiments
Metric SDXL-based Methods SD15-based Methods Ours
infer Time $(s)$ 6 6 9 48 4 3 2 17
[…]
Cross-Modal AdaIN Teacher Model Text Alignment $\uparrow$
0.216
0.223 (+3.2%)
0.228 (+5.5%)
0.235 (+8.7%)
Implementation details. We have implemented our method on top of the latest Adapter-Based Style Transfer approach, named CSGO [37]. To ensure fairness in comparison and mitigate the influence of random initialization, we fixed the initial noise for all methods, as the initial noise can significantly impact the final results [10].
Figure 7: Impact of Style-Based CFG. The proposed Style-Based CFG successfully eliminates unintended style elements such as “green leaves”, “night”, “snow” and “golden leaves”, while text-based CFG fails to address these spurious attributes.
4.1 Comparison with State-of-the-Arts
Qualitative Comparisons. Fig. 6 presents qualitative comparisons with state-of-the-art methods. Existing approaches struggle with prompt alignment due to style overfitting. While CSGO [37], InstantStyle [34], and DEADiff [24] mitigate content leakage, they fail to fully capture prompt-specified details. Conversely, IP-Adapter [38], StyleShot [9] and StyleCrafter [19] exhibit content leakage, leading to images misaligned with the input prompts. In contrast, our method ensures better prompt alignment while maintaining layout stability. More detailed results and explanations could be found in Appendix D.
4.2 Style-Based CFG
4.3 Ablation Study
Teacher Model. In Fig. 8, shows a visualization of cross-attention maps for the word “apple” in the prompt “A red apple” across different model configurations. The first row represents the baseline model, where the attention map fails to effectively focus on the core concept of “apple” from the text prompt. In the second row, the addition of cross-modal AdaIN provides some improvement, but the attention is still diffuse and lacks precise focus. In the third row, with the Teacher Model added, the attention map becomes more concentrated on the target concept, demonstrating a clear improvement in alignment with the prompt. This comparison indicates that it is the Teacher Model that enables the model to better capture and emphasize key elements from the text prompt, resulting in improved attention quality and image generation accuracy. In Fig. 9, this experiment demonstrates the importance of selecting an appropriate denoising timestep to stop the involvement of the Teacher Model. If the Teacher Model’s influence is removed too early, as seen at lower timesteps, issues with layout stability persist, resulting in compositions that lack coherence and include multiple instances of the target object (“apple”). However, if the Teacher Model is involved throughout the full denoising process (e.g., at timestep 50), there is a noticeable loss of style in the generated image. This suggests that while the Teacher Model is crucial for achieving layout stability, an extended involvement can dilute the style elements, making it essential to find a balanced point to discontinue its influence for optimal results. More detailed results and explanations can be found in Appendix C.
Figure 10: More results of our text-driven style transfer model. Illustration of the prompt format used: “A [color] bus”.
Figure 11: The prompts, from left to right, are: “A fluffy white rabbit with pink ears and nose”, “A modern house with a pool.”, “A lovely kitten walking in a garden.”, “A daisy with a ladybug on it”, “A mountain goat on a cliff”, and “A parrot singing a song”.
More Results.  Fig. 10 demonstrates that our method effectively mitigates style overfitting by consistently adapting the target color while preserving structural details. Similarly,  Fig. 11 shows that our model can handle complex prompts without losing style fidelity or scene composition. More detailed results can be found in Appendix E.
5 Conclusion
In conclusion, existing text-driven style transfer faces key issues such as style overfitting and layout instability, which limit the adaptability and coherence of generated images. To address these challenges, we proposed three methods: cross-modal AdaIN for harmonizing style and text features, style-based classifier-free guidance (SCFG) for selective control of stylistic elements, and a Teacher Model to enhance layout stability. Our results confirm that our approach effectively mitigates these issues, improving alignment, stability, and control in style transfer, making it a versatile and robust solution for text-to-image synthesis tasks.
Limitations and Future Work. While our method improves artifact removal and layout stability, the Teacher Model slightly increases inference time. Additionally, generating negative-style image requires expertise and manual effort. Future work could focus on improving efficiency and further exploring strategies to mitigate style overfitting, enabling more adaptive and generalizable style transfer across diverse prompts and visual domains.
Appendix A Implementation Details
We set the random seed to 42 for reproducibility, used 50 inference steps, and applied a uniform guidance scale of 5 across all methods. In the qualitative and quantitative comparison experiments, for the implementation involving the Teacher Model, its participation was specifically limited to the first 20 steps. All experiments were conducted on a single NVIDIA GTX-4090 GPU.